Obtain expression for escape velocity?


1
Expert's answer
2021-01-04T14:39:58-0500

Solution. The escape velocity is this is the speed that must be given to a physical object so that it, moving parallel to the cosmic body, could not fall on it, but at the same time it would remain in a constant orbit. In this case, the object in orbit will be at rest, since two forces will act on it: centrifugal force and gravity.


"\frac{mv^2_{esc}}{R}=\frac{GMm}{R^2}"

where m is body mass; vesc is escape velocity, R is orbital radius; G=6.67x10-11Nm2/kg2 is gravitational constant; M is cosmic body mass. Therefore, the escape velocity is eqaul to


"v_{esc}=\sqrt{\frac{GM}{R}}"

Answer.

"v_{esc}=\sqrt{\frac{GM}{R}}"

where R is orbital radius; G=6.67x10-11Nm2/kg2 is gravitational constant; M is cosmic body mass.